OpenStudy (caitlinnr14):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours. Age Ticket Price Under three years free Three to twelve years $34 Over twelve years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (caitlinnr14):

answer choices (64 – x)34 + (103 – 64) ∙ 45 = C (103 + x)34 + (103 – (64 + x)) ∙ 45 = C (103 - x)34 + (103 – 64) ∙ 45 = C (64 + x)34 + (103 – (64 + x)) ∙ 45 = C

OpenStudy (caitlinnr14):

giv ya a medal!

OpenStudy (texaschic101):

It is the first answer choice :)

OpenStudy (caitlinnr14):

thx! @texaschic101

OpenStudy (texaschic101):

anytime :)